OpenStudy (anonymous):

Find the area of the surface: Part of the sphere x^2+y^2+z^2=4z that lies inside the paraboloid z=x^2+y^2

OpenStudy (anonymous):

@Zarkon @phi @amistre64

OpenStudy (anonymous):

Well your first equation can be rewritten as: \[x^2+y^2+z^2-4z=0 \implies x^2+y^2+(z-2)^2=4\] So a sphere with radius 2 centered at (0,0,2). So you need to find the equation that defined the boundary of the surface. Doing that doesn't seem easy however.

OpenStudy (anonymous):

I'm a bit new to this, so I'm not totally sure how to find the equation of the boundary of the surface. How would I start? Thank you

OpenStudy (anonymous):

Notice that if z=3 for both equations they reduce to: \[x^2+y^2+1=4 \implies x^2+y^2=3; x^2+y^2=z \implies x^2+y^2=3\] This means they meet at: z=3,x^2+y^2=3. So we have something that looks like this: